# Investigating Factuality in Long-Form Text Generation: The Roles of Self-Known and Self-Unknown

### v1

@@ Abstract
Large language models (LLMs) have demonstrated strong capabilities in text understanding and generation. However, they often lack factuality, producing a mixture of true and false information, especially in long-form generation. In this work, we investigates the factuality of long-form text generation across various large language models (LLMs), including GPT-4, Gemini-1.5-Pro, Claude-3-Opus, Llama-3-70B, and Mistral. Our analysis reveals that factuality scores tend to decline in later sentences of the generated text, accompanied by a rise in the number of unsupported claims. Furthermore, we explore the effectiveness of different evaluation settings to assess whether LLMs can accurately judge the correctness of their own outputs: Self-Known (the percentage of supported atomic claims, decomposed from LLM outputs, that the corresponding LLMs judge as correct) and Self-Unknown (the percentage of unsupported atomic claims that the corresponding LLMs judge as incorrect). The results indicate that even advanced models like GPT-4 and Gemini-1.5-Pro fail to achieve perfect Self-Known scores, while their Self-Unknown scores remain notably above zero, reflecting ongoing uncertainty in their self-assessments. Moreover, we find a correlation between higher Self-Known scores and improved factuality, while higher Self-Unknown scores are associated with lower factuality. Interestingly, even without significant changes in the models’ self-judgment (Self-Known and Self-Unknown), the number of unsupported claims can increases, likely as an artifact of long-form generation. These findings show the limitations of current LLMs in long-form generation, and provide valuable insights for improving factuality in long-form text generation.
@@ 1 Introduction
To ensure the reliable use of LLMs, it is also crucial that they possess the ability to recognize not only ”what they know” but also ”what they don’t know.” Recent studies, such as those by Kadavath et al. (2022); Liu et al. (2022); Guerreiro et al. (2023), have shown that language models can assess the validity of their own claims. However, Srivastava et (2023); Yin et al. (2023) have pointed out the limitations of LLMs in recognizing their own knowledge gaps.
In this work, we investigate the factuality patterns of long-form text generation across various LLMs. We first assess the factuality of long-form generation at different relative positions using two annotated datasets and two models: ChatGPT and PerplexityAI (which integrates a search engine). Our findings indicate that sentences generated earlier in the sequence generally demonstrate higher factuality. However, these later-generated sentences also contain more unsupported claims and fewer supported claims.
We further examine the effectiveness of different evaluation settings to assess whether LLMs can accurately evaluate the correctness of atomic claims in their own generated outputs. To quantify the corresponding models’ ability to judge the correctness of atomic claims, we calculate two metrics: the Self-Known score (the percentage of supported atomic claims judged as correct by the LLMs) and the Self-Unknown score (the percentage of unsupported atomic claims judged as incorrect by the LLMs). Our exploration includes three methods, notably a novel approach where the final answer option is replaced with ”None of the above”. This modification appears to provide a more accurate measure of the LLMs’ abilities, as evidenced by a higher flip rate for supported claims and an increasing flip rate at later relative positions. This suggests that the model reassesses its confidence when faced with an option signaling uncertainty. In contrast, the low flip rate for unsupported claims indicates a consistent judgment of their incorrectness. These results suggest a nuanced understanding by LLMs of supported versus unsupported claims and underscore the importance of specific evaluation settings to accurately gauge model performance. Our findings align with human annotations for two LLMs, although some discrepancies, particularly with the PerplexityAI model, suggest gaps in estimation.
@@
1. We explored the factuality patterns of long-form text generation across various model families (GPT-4, Gemini-1.5-Pro, Claude-3-Opus, Llama-3-70B, and Mistral). We found that even the most advanced LLMs typically exhibit lower factuality scores in the later segments of long-form text. Retrieval-Augmented Generation (RAG) systems show a similar trend, although they tend to maintain higher factuality overall.
2. We analyzed Self-Known and Self-Unknown ratios for these LLMs across different segments of their own generated texts. The results showed relatively higher Self-Known scores; however, even strong LLMs (GPT-4, Gemini-1.5-Pro, Claude-3-Opus, etc.) generally achieved only about 50% on the Self-Known score. The Self-Unknown scores were significantly above zero. These findings indicate that even the most advanced LLMs still produce outputs with limited self-acknowledgment ability.
3. We developed a mathematical framework linking Self-Known and Self-Unknown scores to factuality, providing deeper insights into their relationship. Both empirical and theoretical results demonstrate that higher Self-Known scores correspond to improved factuality, while higher Self-Unknown scores are associated with reduced factuality. Moreover, even without significant changes in the models’ self-judgment (Self-Known and Self-Unknown), the number of unsupported claims may still increase, likely reflecting the inherent challenges of long-form generation.
@@ 2.1 Observations
The detailed procedures for computing fractions of different type claims at different relative positions are as following:
1) For every sentence in a generation, we computed the fraction of the number of supported atomic claims, unsupported atomic claims and irrelevant claims; 2) We got the relative position of each sentence, e.g., if it is the third sentence out of six, its relative position would be 3/6 = 50%; 3) We then grouped all sentences that fall within the same relative position range: 0-20%, 20%-40%, 40%-60%, 60%-80% and 80%-100%; 4) Finally, we computed the macro-average percentage within the same relative position group.
Figure 1 (a) shows the ChatGPT results (Figure 8 in the Appendix shows the PerplexityAI results.). We can see that unsupported claims percentage is higher when these sentences are generated later. We hypotheses the possible reasons are the error propagation and these generated claims are with low confidence by LLMs.
We also compute the number of different type claims at different relative positions with the above similar procedures. As shown in Figure 1 (b), LLMs tend to generate less reliable information as they continue the generation. More unsupported claims are included in the continued generation.
@@
Do LLMs recognize that supported claims are indeed supported, and that unsupported claims are unsupported? Do LLMs better identify unsupported claims that appear later in the text compared to those that appear earlier?
@@ 3 Self-Known and Self-Unknown
To investigate these questions, we examine whether the corresponding LLMs recognize their atomic claims by computing two metrics: Self-Known (the percentage of supported atomic claims that the corresponding LLMs judge as correct) and Self-Unknown (the percentage of unsupported atomic claims that the corresponding LLMs judge as incorrect). While there is related work, such as Rajpurkar et al. (2018); Xiong et al. (2024), our approach differs in two key ways: (1) Evaluation is conducted on atomic claims, which are derived from sentences in long-form generation, rather than assigning a score to the entire model output; (2) Our focus is on factuality (whether an atomic claim is true or false), rather than on uncertainty scores (i.e., ”How likely is the above answer to be correct?”).
We explore the computation of Self-Known and Self-Unknown using the following three approaches ( with the corresponding prompt templates provided in Appendix Section A.2):
@@
  • Question-Answering: Given an atomic claim, a question-answer pair can be derived (Trischler et al., 2017; Rajpurkar et al., 2018; Hu et al., 2024) with GPT-4 Turbo. For example, ”Lanny Flaherty is an American.” can be used to derived a question-answer pair (”What nationality is Lanny Flaherty?”, ”American”). Then, given the question and answer, we ask the corresponding LLMs whether the answer is true or false.
  • Question-Answering w/NOA: Similar to the above approach, a question-answer pair is derived according to each atomic claim. One big different is: given question and answer, one more addition choice ( ”None of the above”) (Rajpurkar et al., 2018) is given to the corresponding LLMs. This is a well-defined evaluation because it can check whether the model actually knows the answer of the question, especially if the question is vague or context-information is missing.
@@
In summary, we observed similar results between setting (a) (Direct-Asking) and setting (b) (Question-Answering), and a significant difference between setting (b) (Question-Answering) and setting c (Question-Answering W/ NOA). The deeper analysis between setting b and setting c revealed that setting (c) recognizes atomic claims more confidently and treats atomic claims that flip as unknown. This is why we chose to use setting (c) in the subsequent experiments.
@@ 4 Analysis
We denote the prompt input of LLMs as $\boldsymbol{x}$ and long output of LLMs as $\boldsymbol{y}$. The binary auxiliary label $d=1$ indicates the LLM output is correct and $d=0$ indicates LLM output is wrong.
@@
$=$ $P(d=1,\mathrm{y_{correct}}\mid\boldsymbol{x})\sigma+P(d=1,\mathrm{y_{wrong}}\mid\boldsymbol{x})(1-\sigma)$
$=$ $P(d=1\mid\mathrm{y_{correct}})P(\mathrm{y_{correct}}\mid\boldsymbol{x})\sigma+P(d=1\mid\mathrm{y_{wrong}})P(\mathrm{y_{wrong}}\mid\boldsymbol{x})(1-\sigma)$ (2)
@@
[…]
@@
In Equation 3, we present a method for estimating the factuality score. We use the Self-Known and Self-Unknown results of the corresponding model (ChatGPT) with configuration (c) to estimate the factuality score across different relative positions. As shown in Figure 3, our estimation closely matches the human-annotation results [^3] Due to scarcity of human-annotated, long-form LLM generation datasets, we did not show .
@@ 5 Automatic Results on Additional LLMs
In this section, we examine the trends in factuality, Self-Known scores, and Self-Unknown scores of other advanced LLMs using an automated evaluation tool.
@@ 5.2.1 Experimental Setup
For each LLM, we follow four key steps to obtain experimental results: (1) generating text outputs; (2) filtering the generated content; (3): evaluating factuality; and (4): estimating Self-Known and Self-Unknown scores with the corresponding LLM. For more details on each step, please refer to Section A.5.
@@ 5.2.2 Results
Table 1 presents two results for various LLMs: the average number of atomic claims per generation and the filtered rate. The filtered rate represents the percentage of instances where the LLMs do not provide valuable responses, often due to perceiving insufficient information to generate a meaningful answer. We notice that the behavior of Claude-3-opus and Gemini-1.5-pro is more conservative. These models frequently decide not to provide a valuable response, instead stating something like “I do not have enough verified information”. Figure 4 show results of several powerful LLMs (GPT-4, Gemini-1.5-pro, Claude-3-opus, Llama-3-70B-Instruct, and two Mistral AI models) [^5] Additional LLM results are provided in Figure 9 in the Appendix..
@@
A reasonable question arises: Is the decoding error of LLMs caused by the absence of relevant knowledge? Can Retrieval-Augmented Generation (RAG), which provides additional knowledge, resolve the issue of lower factuality in later stages of generation? In next section, we present some RAG experiments.
@@ 5.3 Retrieval-Augmented Generation
Figure 6(a) demonstrates that the RAG system exhibits significantly lower Self-Known scores and higher Self-Unknown scores. This discrepancy may stem from the corresponding LLM’s lack of prior knowledge regarding the retrieved content in the RAG system, causing it to mistakenly assess accurate information as incorrect.
In these RAG experiments, when all relevant knowledge is incorporated, LLMs show improvements in factuality. However, they still struggle with lower factuality in later generations. This highlights the need for alternative decoding algorithms specifically designed for long-form generation tasks. Implementing more sophisticated decoding techniques could help mitigate the artifacts associated with long-form generation.
@@ 6 Related Work
Recent studies have extensively explored the concepts of Self-Known and Self-Unknown in language models. For instance, Kadavath et al. (2022); Liu et al. (2022); Guerreiro et al. (2023) demonstrated that language models are capable of assessing the validity of their own claims and predicting their ability with answering true/false questions accurately. Meanwhile, Srivastava et (2023); Yin et al. (2023) highlighted the limitations of LLMs in acknowledging their unknowns, focusing on their ability to recognize unknown knowledge. In our work, we specifically investigate whether LLMs can identify and reconsider unsupported claims generated from their own outputs. Our results indicate that LLMs struggle to accurately judge unsupported atomic claims from their own generations. We find that a lower Self-Unknown score or a higher Self-Known score corresponds to higher factuality.
@@ 7 Conclusion
In this study, we investigate the factuality of long-form text generation across different model families and relative positions. Our findings reveal a trend of lower factuality in sentences generated later in the sequence. Additionally, we propose methods for enabling LLMs to accurately assess the correctness of atomic claims derived from their own outputs. We introduce an estimation of factuality using Self-Known and Self-Unknown scores, finding that higher Self-Known scores correlate with increased factuality, whereas higher Self-Unknown scores correlate with decreased factuality. These insights provide valuable directions for enhancing factuality in text generation.
Appendix A Appendix
A.1 LLMs

### v2

@@ Abstract
Large language models (LLMs) have demonstrated strong capabilities in text understanding and generation. However, they often lack factuality, producing a mixture of true and false information, especially in long-form generation. In this work, we investigates the factuality of long-form text generation across various large language models (LLMs), including GPT-4, Gemini-1.5-Pro, Claude-3-Opus, Llama-3-70B, and Mistral. Our analysis reveals that factuality tend to decline in later sentences of the generated text, accompanied by a rise in the number of unsupported claims. Furthermore, we explore the effectiveness of different evaluation settings to assess whether LLMs can accurately judge the correctness of their own outputs: Self-Known (the percentage of supported atomic claims, decomposed from LLM outputs, that the corresponding LLMs judge as correct) and Self-Unknown (the percentage of unsupported atomic claims that the corresponding LLMs judge as incorrect). Empirically, we observe a positive correlation between higher Self-Known scores and improved factuality, whereas higher Self-Unknown scores are associated with reduced factuality. Interestingly, the number of unsupported claims can increase even without significant changes in a model’s self-judgment scores (Self-Known and Self-Unknown), likely as a byproduct of long-form text generation. We also derive a mathematical framework linking Self-Known and Self-Unknown scores to factuality: $\textrm{Factuality}=\frac{1-\textrm{Self-Unknown}}{2-\textrm{Self-Unknown}-\textrm{Self-Known}}$, which aligns with our empirical observations. Additional Retrieval-Augmented Generation (RAG) experiments further highlight the limitations of current LLMs in long-form generation and underscore the need for continued research to improve factuality in long-form text.
Investigating Factuality in Long-Form Text Generation: The Roles of Self-Known and Self-Unknown
Lifu Tu, Rui Mengthanks: Now at Google., Shafiq Joty, Yingbo Zhou, Semih Yavuz Salesforce AI Research
@@ 1 Introduction
To ensure the reliable use of LLMs, it is also crucial that they possess the ability to recognize not only "what they know" but also "what they don’t know." Recent studies, such as those by Kadavath et al. (2022); Liu et al. (2022); Guerreiro et al. (2023), have shown that language models can assess the validity of their own claims. However, Srivastava et (2023); Yin et al. (2023) have pointed out the limitations of LLMs in recognizing their own knowledge gaps.
In this work, we investigate the factuality of long-form text generation across various LLMs. We first check the factuality of long-form generation at different relative positions using two annotated datasets and two models: ChatGPT and PerplexityAI (which integrates a search engine). Our findings verify that sentences generated earlier in the sequence generally demonstrate higher factuality. However, these later-generated sentences contain more unsupported claims and fewer supported claims.
To explain this phenomenon, we try to examine whether LLMs exhibit reduced knowledge in later generations with wo metrics: the Self-Known score (the percentage of supported atomic claims judged as correct by the LLMs) and the Self-Unknown score (the percentage of unsupported atomic claims judged as incorrect by the LLMs). These two metrics are used to quantify the corresponding models’ ability to judge the correctness of atomic claims. In order to accurately compute the two metrics, we have tried three different approaches, one of which is a novel approach that adds an answer option: ‘None of the above’. This modification appears to provide a more accurate measure of the LLMs’ abilities, as evidenced by a higher flip rate for supported claims and an increasing flip rate at later relative positions. This suggests that the model reassesses its confidence when faced with an option signaling uncertainty. In contrast, the low flip rate for unsupported claims indicates a consistent judgment of their incorrectness. These results suggest a nuanced understanding by LLMs of supported versus unsupported claims and underscore the importance of specific evaluation settings to accurately gauge model performance. Our findings align with human annotations for two LLMs, although some discrepancies, particularly with the PerplexityAI model, suggest gaps in estimation.
Later, we apply this modified approach to compute the Self-Known and Self-Unknown scores across various LLMs, including GPT-4, Gemini-1.5-Pro, Claude-3-Opus, Llama-3-70B, and Mistral. We also develop a mathematical framework that links these scores to factuality. Overall, both empirical and theoretical results demonstrate a strong relationship between factuality and the Self-Known and Self-Unknown scores. We observe that these scores can vary significantly across different models. However, even when the Self-Known and Self-Unknown scores remain relatively stable, the number of unsupported claims tends to increase in later parts of the generated text. This suggests that lower factuality in later sentences is not solely due to score changes, but also influenced by error propagation and diminished model knowledge over time.
@@
1. We explore the factuality patterns of long-form text generation across various model families (GPT-4, Gemini-1.5-Pro, Claude-3-Opus, Llama-3-70B, and Mistral). We find that even the most advanced LLMs typically exhibit lower factuality scores in the later segments of long-form text. Retrieval-Augmented Generation (RAG) systems show a similar trend, although they tend to maintain higher factuality overall.
2. We analyze Self-Known and Self-Unknown ratios across different segments of generated text. While Self-Known scores are relatively high, even the strongest LLMs (e.g., GPT-4, Gemini-1.5-Pro, Claude-3-Opus) average only around 50%, with Self-Unknown scores remaining well above zero. This suggests that even advanced models struggle to recognize the limits of their own knowledge.
3. We develop a mathematical framework linking Self-Known and Self-Unknown scores to factuality. Empirical and theoretical results show higher Self-Known scores improve factuality, while higher Self-Unknown scores reduce it. Notably, unsupported claims can increase even without major changes in self-judgment, highlighting challenges in long-form generation.
4. We find that Retrieval-Augmented Generation (RAG), which supplies needed knowledge, can improve factuality. However, it fails to fully address the issue of lower factuality at a later position. This highlights the need for alternative framework specifically designed for long-form generation tasks.
@@ 2.1 Observations
The detailed procedures for computing fractions of different type claims at different relative positions are as following: 1) Calculate the fraction of supported, unsupported, and irrelevant claims for each sentence; 2) Determine each sentence’s relative position in the generation , e.g., if it is the third sentence out of six, its relative position would be 3/6 = 50%; 3) Group sentences into relative position ranges: 0-20%, 20%-40%,, etc.; 4) Compute the macro-average claim percentages within each group
Figure 1 (a) shows ChatGPT results (PerplexityAI results are in the Appendix). Unsupported claims increase in later-generated sentences. Figure 1 (b) further shows that as generation continues, LLMs produce more unreliable and unsupported claims.
@@
Is the phenomenon above due to LLMs having less knowledge about later generations? Can LLMs recognize when claims are supported and when they are not? Do LLMs identify unsupported claims more effectively when they appear later in the text compared to earlier ones?
@@ 3 Self-Known and Self-Unknown
To investigate these questions, we examine whether the corresponding LLMs recognize their atomic claims by computing two metrics: Self-Known (the percentage of supported atomic claims that the corresponding LLMs judge as correct) and Self-Unknown (the percentage of unsupported atomic claims that the corresponding LLMs judge as incorrect). While there is related work, such as Rajpurkar et al. (2018); Xiong et al. (2024), our approach differs in two key ways: (1) Evaluation is conducted on atomic claims, which are derived from sentences in long-form generation, rather than assigning a score to the entire model output; (2) Our focus is on factuality (whether an atomic claim is true or false), rather than on uncertainty scores (i.e., "How likely is the above answer to be correct?").
We explore the computation of Self-Known and Self-Unknown using the following three approaches ( with the corresponding prompt templates provided in Appendix Section B):
@@
  • Question-Answering: Given an atomic claim, a question-answer pair can be derived (Trischler et al., 2017; Rajpurkar et al., 2018; Hu et al., 2024) with GPT-4 Turbo. For example, "Lanny Flaherty is an American." can be used to derived a question-answer pair ("What nationality is Lanny Flaherty?", "American"). Then, given the question and answer, we ask the corresponding LLMs whether the answer is true or false.
  • Question-Answering w/NOA: Similar to the above approach, a question-answer pair is derived according to each atomic claim. One big different is: given question and answer, one more addition choice ( "None of the above") (Rajpurkar et al., 2018) is given to the corresponding LLMs. This is a well-defined evaluation because it can check whether the model actually knows the answer of the question, especially if the question is vague or context-information is missing.
@@
In summary, we observed similar results between setting (a) (Direct-Asking) and setting (b) (Question-Answering), and a significant difference between setting (b) (Question-Answering) and setting c (Question-Answering W/ NOA). The deeper analysis between setting (b) and setting (c) revealed that setting (c) recognizes atomic claims more confidently and treats atomic claims that flip as unknown. This is why we chose to use setting (c) in the subsequent experiments.
@@ 4 Analysis
We denote the prompt input of LLMs as $\boldsymbol{x}$ and long output of LLMs as $\boldsymbol{y}$. The binary auxiliary label $d=1$ indicates the LLM output is factual correct and $d=0$ indicates LLM output is wrong.
@@
$=$ $P(d=1,\mathrm{y_{correct}}\mid\boldsymbol{x})*\sigma+$
$P(d=1,\mathrm{y_{wrong}}\mid\boldsymbol{x})*(1-\sigma)$
$=$ $P(d=1\mid\mathrm{y_{correct}})*P(\mathrm{y_{correct}}\mid\boldsymbol{x})*\sigma+$
$P(d=1\mid\mathrm{y_{wrong}})*P(\mathrm{y_{wrong}}\mid\boldsymbol{x})*(1-\sigma)$ (2)
@@
Given $\textrm{Self-Unknown}\in[0,1]$ and $\textrm{Self-Known}\in[0,1]$, the factuality score increases when the Self-Known score is increased or the Self-Unknown score is decreased. This matches our observations in Section 3 and Figure 2 (c).
@@
In Equation 3, we present a method for estimating the factuality score. We use the Self-Known and Self-Unknown results of the corresponding model (ChatGPT) with configuration (c) to estimate the factuality score across different relative positions. As shown in Figure 3, our estimation closely matches the human-annotation results.
@@ 5 Automatic Results on Additional LLMs
In the previous section, our experiments were conducted using human-annotated factuality data. In this section, we first introduce an automated tool for factuality evaluation. Then, using the proposed approach from Section 3 to compute Self-Known and Self-Unknown scores, we analyze the trends in factuality, Self-Known, and Self-Unknown scores across other advanced LLMs
@@ 5.2.1 Experimental Setup
For each LLM, we follow four key steps to obtain experimental results: (1) generating text outputs; (2) filtering the generated content; (3): evaluating factuality; and (4): estimating Self-Known and Self-Unknown scores with the corresponding LLM. For more details on each step, please refer to Appendix Section E.
@@ 5.2.2 Results
Figure 4 show results of several powerful LLMs ( Gemini-1.5-pro, Claude-3-opus, and two Mistral AI models). Two additional LLMs (GPT-4, and Llama-3-70B-Instruct) results are provided in Figure 9 in the Appendix.
@@
Some reasonable questions arise: Are decoding errors in LLMs caused by a lack of relevant knowledge? Can Retrieval-Augmented Generation (RAG), which supplies additional context, address the decline in factuality during later stages of generation? In the next section, we present our exploration of RAG-based experiments across different LLMs.
@@ 5.3 Retrieval-Augmented Generation
Figure 6 (b) demonstrates that the RAG system exhibits significantly lower Self-Known scores and higher Self-Unknown scores. This discrepancy may stem from the corresponding LLM’s lack of prior knowledge regarding the retrieved content in the RAG system, causing it to mistakenly assess accurate information as incorrect.
In these RAG experiments, incorporating all relevant knowledge leads to improved factuality in LLMs. However, they still exhibit a decline in factuality during later stages of generation. This underscores the need for alternative frameworks specifically tailored to long-form generation tasks. For instance, employing more sophisticated decoding strategies may help mitigate the challenges associated with long-form generation.
@@ 6 Related Work
Recent studies have extensively explored the concepts of Self-Known and Self-Unknown in language models. For instance, Kadavath et al. (2022); Liu et al. (2022); Guerreiro et al. (2023) demonstrated that language models are capable of assessing the validity of their own claims and predicting their ability with answering true/false questions accurately. Meanwhile, Srivastava et (2023); Yin et al. (2023) highlighted the limitations of LLMs in acknowledging their unknowns, focusing on their ability to recognize unknown knowledge. In our work, we specifically investigate whether LLMs can identify and reconsider unsupported claims generated from their own outputs. Our results indicate that LLMs struggle to accurately judge unsupported atomic claims from their own generations. We also find that a lower Self-Unknown score or a higher Self-Known score corresponds to higher factuality.
@@ 7 Conclusion
n this study, we investigate the factuality of long-form text generation across different language model families and at various stages of generation. We observe a consistent decline in factuality in sentences generated later in the sequence. To understand the underlying causes, we explore two possible factors: diminished self-knowledge in later generations and the accumulation of earlier generation errors (i.e., error propagation). To analyze this, we introduce the concepts of Self-Known and Self-Unknown scores, which measure a model’s confidence in its own knowledge. We find that current LLMs struggle to maintain factual accuracy over extended generations, partly due to limitations in their internal knowledge representation and propagation mechanisms. Addressing these challenges requires further research. Promising directions include the development of external factuality evaluation modules (e.g., dedicated judge models) and the design of more robust decoding strategies tailored to long-form generation
8 Limitations
Following are limitations in our work.
Evaluation of Self-Know and Self-Unknown
In this work, we design three different methods for estimating Self-Known and Self-Unknown scores on LLMs’ own generation. We find that the third setting (c), which includes the option "None of the above," is effective in determining whether LLMs can accurately judge the correctness of claims generated from their own outputs. Although our results show that these scores are well aligned with the estimation of factuality scores using Equation 3, exploring better methods for evaluating the correctness of claims with LLMs would still be beneficial for future study.
Factuality Evaluation
In this work, we limit the domain of long-form generation to ensure accurate factuality evaluation. The concern is that broadening the topic range might compromise the accuracy of our factuality assessments, rendering our study less effective. Recently, evaluation tools (Guan et al., 2024; Es et al., 2023; Tang et al., 2024; Wei et al., 2024) have been explored. With stronger tools, it is possible to explore a wider range of domains beyond Wikipedia.
Moreover, in this work, we primarily focus on factuality precision. However, considering factuality recall is also important, as it ensures that the omission of significant pieces of information is penalized during evaluation. By incorporating both precision and recall, we can achieve a more comprehensive and accurate assessment of factuality in long-form generation.
Appendix A LLMs